Minimize the following multi level function, F = ABC + [A + D][A' + C']

a) F=(AD)'+AB'C

b) F=AD'+A(BC)'

c) F=(AD)'+(AB)'

d) F=AD+(AB)'C
Answer #1

Fe ABC+ (6 +22 Alts) Fa ABC+ AA + Actad tco CAA=o) F- A B C + Actadtce + ( xy +2 = (2+2) (ytz) = A (B C+ c) + Aldted f = A